OpenStudy (anonymous):

Solve the following equation. One point is given for showing steps and one point given for the correct solution. (x - 3) /x = (x - 3) / (x - 6)

OpenStudy (shubhamsrg):

any attempt ?

OpenStudy (anonymous):

I dont understand how to do it..

OpenStudy (anonymous):

can you help me?

OpenStudy (whpalmer4):

\[\frac{(x-3)}{x} = \frac{(x-3)}{(x-6)}\] Whenever you have something like that, cross-multiplying is a good place to start.

OpenStudy (anonymous):

Okay thank you!

OpenStudy (anonymous):

you might could eyeball it as well

OpenStudy (anonymous):

for example, if \(x=3\) then both the left and right sides are zero and if \(x\neq 3\) then since the numerators are equal, the denominators would have to be equal as well. but there is not much chance for \(x=x-6\)